The Shirt Shop had the following transactions for T-shirts for 2018, its first year of operations:

Jan. 20 Purchased 440 units @ $ 9 = $ 3,960
Apr. 21 Purchased 150 units @ $ 10 = 1,500
July 25 Purchased 240 units @ $ 11 = 2,640
Sept. 19 Purchased 80 units @ $ 12 = 960

During the year, The Shirt Shop sold 720 T-shirts for $17 each.


a. Compute the amount of ending inventory The Shirt Shop would report on the balance sheet, assuming the following cost flow assumptions: (1) FIFO, (2) LIFO, and (3) weighted average. (Round intermediate calculations to 2 decimal places and final answers to nearest whole dollar amount.)
Answer #1
Units Unit cost Total
Jan. 20 440 9 3960
Apr. 21 150 10 1500
July 25 240 11 2640
Sept. 19 80 12 960
Total 910 9060
Ending inventory units 190 =910-720
Average cost 9.96
a
Amount of ending inventory :
FIFO 2170 =(80*12)+(190-80)*11
LIFO 1710 =190*9
Weighted average 1892 =190*9.96